Circular No. 6312

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

PROBABLE GRAVITATIONAL MICROLENSING EVENT The MACHO collaboration (C. Alcock, T. Axelrod, R. Allsman, D.

Alves, A. Becker, D. Bennett, K. Cook, K. Freeman, K. Griest, J. Guern, M. Lehner, S. Marshall, B. Peterson, M. Pratt, P. Quinn, A. Rodgers, C. Stubbs, W. Sutherland, and D. Welch; Lawrence Livermore National Lab., Center for Particle Astrophysics, and Mount Stromlo Observatory) reports the discovery of a likely gravitational microlensing event in progress toward the Large Magellanic Cloud: “The star in question is located at α = 5^{h}24^{m}03^{s}.2, δ = −68^{o}49′12′′ (equinox 2000.0). The star was constant at R = 20.9, V = 21.1 during 1993–1995 and brightened by 0.9 mag during the month prior to 1996 Feb. 14. The color appears unchanged. The current microlensing fit predicts that the star should reach a peak of V < 19.9 between Feb. 22 and Mar. 31 and then decline symmetrically. Accurate photometry and low-resolution spectroscopy of this star at regular intervals over the next few weeks will be valuable to test the microlensing interpretation. Photometric measurements may possibly detect lightcurve deviations that are due to a breakdown of the usual point-source, pointlens, or constant-velocity assumptions. If these exotic effects can be detected, it is usually possible to estimate the distance to the lensing object. A finding chart and additional information are available by anonymous

FTP at darkstar.astro.washington.edu:/macho/Alerts/96-LMC-1 or
on the WWW at http://darkstar.astro.washington.edu/. Potential

observers are requested to contact macho@astro.washington.edu, D. Bennett at telephone 510-423-0656, or S. Marshall at 510-422-4872, to coordinate observations.”

COMET C/1995 O1 (HALE-BOPP) M. C. Senay, Institute for Astronomy (IfA), University of Hawaii; H. E.

Matthews, Joint Astronomy Centre, Hilo; and D. Jewitt, IfA, write: “We observed C/1995 O1 using the 15-m James Clerk Maxwell Telescope on Feb. 10 UT. The 345-GHz J = 3-2 transition of carbon monoxide was detected with peak antenna temperature 0.22 K, area 0.09 ± 0.01 K km/s, and central velocity −33.06 ± 0.02 km/s. The preliminary CO production rate computed using coma temperature(s) 10 K (50 K) is 3.7 (1.2) × 10^{28}, corresponding to the mass-loss rate(s) 1700 (560) kg/s. Continuous CO outgassing has now been observed for 5 months (cf. IAUC 6234). We estimate the total loss to date at 2 × 10^{35} molecules (1 × 10^{10} kg).”

1996 February 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green